The mean weight of luggage checked by a randomly selected economic class passenger flying between two cities on a certain airline is 40Ib, and the standard deviation is 10 lb. The mean and standard deviation for a business class passenger are 30lb and 6 lb respectively. If there are 50 economic class and 12 business class passengers on the flight, what is the variance of total luggage weight?

Select one:  a. 2360  b. 5000  c. 432  d. 5432 


1
Expert's answer
2021-11-01T16:11:22-0400

Mean luggage weight of tourist class passenger is 40 LB and standard deviation is 10 LB.

Mean luggage weight of business class is 30 LB and standard deviation is 6 LB.

Therefore, mean of 12 business class and 50 economic class passenger is "(50 \times 40+12 \times 30)=560"

"variance = 50 \times 10^2+12 \times 6^2=5432"

Answer: d. 5432